# Multi-agent cooperation through learning-aware policy gradients

### v1

@@ 4.1 The iterated prisoner’s dilemma
We model each agent through a tabular policy $\pi^{i}(a_{t}^{i}\mid x_{t}^{i};\phi^{i})$ that depends only on the previous action of both agents, $x_{t}^{i}=(a^{1}_{t-1},a^{2}_{t-1})$. Their behavior is thus fully specified by five parameters, which determine the probability of cooperating in response to the four possible previous action combinations together with the initial cooperation probability. For this game, the discounted expected return $J^{i}(\phi^{1},\phi^{2})$ can be calculated analytically. We exploit this property and optimize policies by performing exact gradient ascent on the expected return, with gradients computed by automatic differentiation (c.f. Appendix B for details). Results can be found in Fig. 3.
@@ 4.2 Explaining cooperation through learning awareness
Finding 1: Learning-aware agents extort naive learners. We first pit naive against learning-aware agents. We find that the latter develop extortion policies which force naive learners onto unfair cooperation, in the spirit of the zero-determinant extortion strategies discovered by Press and Dyson (2012). Even when a learning-aware agent is initialized at pure defection, maximizing the shaping objective of Eq. 3 lets it escape mutual defection (see Fig. 3A).
@@
Finding 3: Cooperation emerges within groups of naive and learning-aware agents. Findings 1. and 2. motivate studying learning in a group containing both naive and learning-aware agents, with every agent in the group trained against each other. This mixed group setting yields a sum of two distinct shaping objectives, which depend on whether the agent being shaped is learning-aware or naive. The gradients resulting from playing against naive learners pull against mutual defection and towards extortion, while those resulting from playing against other learning-aware agents push away from extortion towards cooperation. Balancing these competing forces leads to robust cooperation, see Fig. 3C (left). By contrast, a pure group of learning-aware agents cannot escape mutual defection, see Fig. 3C (right). This can be explained by the fact that the agents can no longer observe other players learning, and must deal again with a non-stationary problem. The resulting gradients do not therefore contain information on the effects of unconditional defection on the future strategies of co-players, or that policies in the vein of tit-of-tat can shape co-players towards more cooperation.
@@
5 Revisiting Learning with Opponent-Learning Awareness
We now leverage our results from Sections 3-4 to explain how and when cooperation arises in Learning with Opponent-Learning Awareness (LOLA; Foerster et al. (2018a)), the seminal work that spearheaded the learning awareness field. Furthermore, we will use COALA-PG to derive a new LOLA gradient estimator that does not require higher-order derivative estimates.
LOLA considers a POSG (c.f. Section 2.2) with agent policies $(\phi^{i},\phi^{-i})$, and expected return $J^{i}(\phi^{i},\phi^{-i})$. Instead of estimating the naive gradients $\nabla_{\phi^{i}}J^{i}(\phi^{i},\phi^{-i})$, LOLA anticipates that the co-players update their parameters with $M$ naive gradient steps, and estimates the total gradient through this look-ahead objective:
@@ 5 Revisiting Learning with Opponent-Learning Awareness
with $\frac{\mathrm{d}}{\mathrm{d}\phi^{i}}$ the total derivative taking into account the effect of $\phi^{i}$ on the parameter updates $\Delta_{q}\phi^{-i}$, and $\frac{\partial}{\partial\phi^{-i}}$ the partial derivative. Note that Eq. 5 considers the LOLA-DICE update (Foerster et al., 2018b), an improved version of the LOLA update. In reality, all players update their parameters with the LOLA-DICE update equation 5, hence the look-ahead assumption with naive gradients is inconsistent, sometimes hampering the performance of LOLA (Willi et al., 2022; Letcher et al., 2019).
5.1 Explaining when and how cooperation arises with LOLA
@@
In Section 4, we showed that the two main ingredients for learning to cooperate with selfish objectives are (i) observe that one’s actions influence the future behavior of others, providing shaping gradients pulling away from defection towards extortion, and (ii), also play against other extortion agents immune to being shaped on the fast timescale, providing gradients pulling away from extortion towards cooperation. We then showed that both ingredients can be combined by training agents in a heterogeneous group containing both naive and learning-aware agents.
We can explain the emerging cooperation in LOLA by observing that LOLA also combines both ingredients, albeit in a different manner from the heterogeneous group setting. By considering the look-ahead objective equation 5, LOLA computes gradients to shape naive learners performing $M$ naive gradient steps. Unique to LOLA however, the naive learners in the look-ahead objective equation 5 are initialized with the parameters of the other LOLA agents $\phi^{-i}$. If only few look-ahead steps are used, the updated naive learner parameters stay close to $\phi^{-i}$, mimicking playing against other extortion agents resulting in emergent cooperative behavior.
Figure 4A shows that indeed, LOLA-DICE with few look-ahead steps leads to cooperation on the iterated prisoner’s dilemma. However, when more look-ahead steps are used, the naive learner moves too far away from its $\phi^{-i}$ initialization, removing the second ingredient, leading to defection. In figure 4, we take the policy resulting from LOLA-training with many look-ahead steps, and train a new randomly initialized naive learner against this fixed LOLA policy. The results show that the LOLA policy extorts the naive learner into unfair cooperation, confirming that with many look-ahead steps, only a shaping incentive is present in the LOLA update, resulting in extortion policies. Hence, the low reward in Fig. 4A for LOLA agents with many look-ahead steps does not result from unconditional defection, but instead both LOLA policies trying to extort the other one. Finally, we can improve the performance of LOLA with many look-ahead steps by explicitly introducing ingredient 2 through adding the partial gradient $\frac{\partial}{\partial\phi^{-i}}J^{i}(\phi,\phi^{-i})$ to the LOLA update equation 5 (c.f. Fig. 4C).
@@
Figure 4: (A) Performance of two agents trained by LOLA on the iterated prisoner’s dilemma with analytical gradients for various look-ahead steps (only the performance of the first agent is shown). (B) Performance of a randomly initialized naive learner trained against the fixed LOLA 20 look-aheads policy taken from the end of training of (A). (C) Same setting as (A), but with the naive gradient $\lambda\frac{\partial}{\partial\phi^{-i}}J^{i}(\phi,\phi^{-i})$ added to the LOLA update, with $\lambda$ a hyperparameter (c.f. Appendix B). Shaded regions indicate standard error computed over 64 seeds.
5.2 A higher-derivative-free LOLA gradient estimator
Despite the apparent dissimilarities of the two algorithms, we show in Theorem 5.1 that as a special case of COALA-PG, we can estimate the gradient of a similar objective to the LOLA-DICE method. We consider a mixed agent group of meta agents $(\phi^{i},\phi^{-i})$, and naive learners.
Theorem 5.1.
@@ 6 Experimental analysis of policy gradient implementations
The results presented on the previous section were obtained by performing gradient ascent on analytical expected returns. This assumes knowledge of co-player parameters, and it is only possible on a restricted number of games which admit closed-form value functions. We now move to the general reinforcement learning setting, with the main aim of studying the performance of the COALA-PG learning rule. We consider both the iterated prisoner’s dilemma and a simplified version of the CleanUp game, a challenging general-sum game belonging to the Melting Pot suite of environments (Agapiou et al., 2023). We compare to the M-FOS meta reinforcement learning algorithm (Lu et al., 2022), which we take as a best-in-class baseline method, as well as to an ablation motivated by our batched POMDP.
@@ 6.1 Agents trained with COALA-PG master the iterated prisoner’s dilemma
We train a long-context sequence policy $\pi^{i}(a^{i,b}_{l}\mid h_{l}^{i,b};\phi^{i})$ with the COALA-PG rule to play the (finite) iterated prisoner’s dilemma, see Appendix A. We choose a Hawk recurrent neural network as the policy backbone (De et al., 2024). Hawk models achieve transformer-level performance at scale, but with time and memory costs that grow only linearly with sequence length. This allows processing efficiently the long history context $h_{l}^{i,b}$, which contains all actions played by the agents across episodes, from their first round up to current round $l$. Based on the results of the preceding section, we consider a mixed group setting, pitting COALA-PG-trained agents against naive learners as well as other equally capable learning-aware agents. Naive learners are equipped with the same policy architecture as the agents trained by COALA-PG, but their context is limited to the current inner game history $x_{t}^{i,b}$.
@@
Figure 5: Agents trained by COALA-PG play iterated prisoner’s dilemma. Left: When trained against naive agents only, COALA-PG-trained agents extort the latter and reach considerably higher reward than other baseline agents. The stars ($\star$) indicate overlapping curves of the corresponding color at that point Center: When analyzing the behavior of the agents within one meta-episode, we observe COALA-PG-trained agents shaping naive co-players, leading to low defection rate in the beginning, which is then exploited towards the end. M-FOS on the other hand defects from the beginning, achieving lower reward, thus failing to properly optimize the shaping problem. Batch-unaware COALA-PG performs identically to M-FOS and is therefore omitted. Right: Average performance of meta agents playing against other meta agents, when training a group of meta agents against a mixture of naive and other meta agents. Such agents trained with COALA-PG cooperate when playing against each other, but fail to do so when trained with baseline methods. When removing naive agents from the pool, meta agents also fail to cooperate, as predicted in Section 3. Shaded regions indicate standard deviation computed over 5 seeds.
In Fig. 5, we see that COALA-PG reproduces the analytical game findings reported in the previous section: learning-aware agents cooperate with other learning-aware agents, and extort naive learners. Importantly, the identity of each agent is not revealed to a learning-aware agent, which must therefore infer in-context the strategy used by the player it faces.
@@
By contrast, the whole group falls into defection when training the exact same sequence model with the M-FOS rule, which weighs disproportionately future vs. current episode returns. We note that the experiments reported by Lu et al. (2022) were performed with a tabular policy and analytical inner game returns, for which cooperation could be achieved with the M-FOS rule. This shows how crucial the unbiased policy gradient property of COALA-PG is for co-player shaping by meta reinforcement learning to succeed in practice. The same failure to beat defection occurs when using a naive policy gradient meta-update ablation which does not take co-player batching into account. We refer to Appendix E for expressions for this baseline as well as the M-FOS rule.
@@ 6.2 Agents trained with COALA-PG cooperate on a sequential social dilemma
Figure 6: Agents trained by COALA-PG against naive agents only successfully shape them in CleanUp. Left: COALA-PG-trained agents better shape naive opponents compared to baselines, obtaining higher return. Center left & center right & right: Analyzing behavior within a single meta-episode after training reveals that COALA outperforms baselines and shapes naive agents, (i) collecting overall higher reward across the episode, (ii) exhibiting an improved cleaning-harvesting ratio, and (iii) being less often zapped. Shaded regions indicate standard deviation computed over 5 seeds.
@@
Finally, we consider CleanUp-lite, a simplified two-player version of the CleanUp game, which is part of the Melting Pot suite of multi-agent environments (Agapiou et al., 2023). We briefly describe the game here, and provide additional details on Appendix A. On a high level, CleanUp-lite is a two-player game that models the social dilemma known as the tragedy of the commons (Hardin, 1968). A player receives rewards by picking up apples. Apples are spontaneously generated in an orchard, but the rate of generation is inversely proportional to the pollution level of a nearby river. Agents can spend time cleaning the river to reduce the pollution level, and thus increase the apple generation rate. In a single-player game, an agent would balance out cleaning and harvesting to maximize the return. In a multi-player setting however, this gives room for a ‘freerider’ who never cleans and always harvests, letting the opponent clean instead. At any time point, agents can ‘zap’ the opponent, which would result in the opponent being frozen for a number of time steps, unable to harvest or clean. In contrast to matrix games where cooperation corresponds to a single action, this game is a sequential social dilemma (Leibo et al., 2017), where cooperation involves orchestrating sequences of actions.
@@
As in the previous section, we model agent behavior through Hawk sequence policies, and compare COALA-PG to the same baseline methods as before. Naive agents here would learn too slowly if initialized from scratch, and are therefore handled differently, see Appendix A.2.2. In Figs. 6 and 7, we see that agents trained by the COALA-PG rule reach significantly higher returns than all baselines, establishing a mutual cooperation protocol with other learning-aware agents, while exploiting naive ones. We further describe below the qualitative behavior found in the simulations.
Exploitation of naive agents.
@@
COALA-PG-trained agents shape the behavior of naive ones to their advantage. Our behavioral analysis reveals three salient features. First, COALA-PG successfully shapes naive opponents to zap less often throughout the meta-episode. Less overall zapping means that agents can harvest more apples while the pollution level is low, thus increasing the overall reward. Second, when COALA-PG shapes naive learners, overall, the pollution level is lower on average. Looking closer, both baselines spend all of their time harvesting, delegating the cleaning entirely to the naive agents. In that case, the naive agent quickly alternates between cleaning and picking up crumbs in the orchard, but it is not incentivized to bring the pollution down significantly. On the other hand, COALA-PG-trained agents also participate in cleaning (albeit less than naive ones), which brings the pollution level further down, thus increasing the apple production. Third, the increased availability in apples in the learning-aware vs. naive setting is disproportionately picked up by the learning-aware agent, who increases its reward w.r.t. baselines by about 60%, as opposed to about 40% for the naive agent, who therefore also benefits from the increased apples.
@@
Figure 7: Agents trained with COALA-PG against a mixture of naive and other meta agents learn to cooperate in CleanUp-lite. Left: COALA-PG-trained agents obtain higher reward than baseline agents. Center & right: COALA-PG leads to higher apple production because of lower pollution through more cleaning and lower zapping rates. Shaded regions indicate standard deviation computed over 5 seeds.
Learning-aware agents cooperate.
@@
We see similar trends when introducing other COALA-PG-trained agents in the game, see Fig. 7. Essentially, COALA-PG allows for higher apple production because of lower pollution, and lower zapping rate. We see that over training time, cleaning rate goes up, and zapping rate goes down. Interestingly, the cleaning (38 % vs 17 %) and zapping rates (5 % vs 7%) averaged over the meta-episode are higher resp. lower than in the pure shaping setting (i.e., with naive agents only), indicating that learning-aware agents mutually shape each other to clean more and zap less.

### v2

@@ 4.1 The iterated prisoner’s dilemma
We model each agent through a tabular policy $\pi^{i}(a_{t}^{i}\mid x_{t}^{i};\phi^{i})$ that depends only on the previous action of both agents, $x_{t}^{i}=(a^{1}_{t-1},a^{2}_{t-1})$. Their behavior is thus fully specified by five parameters, which determine the probability of cooperating in response to the four possible previous action combinations together with the initial cooperation probability. For this game, the discounted expected return $J^{i}(\phi^{1},\phi^{2})$ can be calculated analytically. We exploit this property and optimize policies by performing exact gradient ascent on the expected return (c.f. Appendix C for details).
@@ 4.2 Explaining cooperation through learning awareness
Finding 1: Learning-aware agents extort naive learners. We first pit naive against learning-aware agents. We find that the latter develop extortion policies which force naive learners onto unfair cooperation, similar to the zero-determinant extortion strategies discovered by Press and Dyson (2012) (c.f. Appendix C.1). Even when a learning-aware agent is initialized at pure defection, maximizing the shaping objective of Eq. 2 lets it escape mutual defection (see Fig. 3A).
@@
Finding 3: Cooperation emerges within groups of naive and learning-aware agents. Findings 1. and 2. motivate studying learning in a group containing both naive and learning-aware agents, with every agent in the group trained against each other. This mixed group setting yields a sum of two distinct shaping objectives, which depend on whether the agent being shaped is learning-aware or naive. The gradients resulting from playing against naive learners pull against mutual defection and towards extortion, while those resulting from playing against other learning-aware agents push away from extortion towards cooperation. Balancing these competing forces leads to robust cooperation, see Fig. 3C (left). Intriguingly, mutual unconditional defection is no longer a Nash equilibrium in this mixed group setting, and agents initialized with unconditional defection policies learn to cooperate (see Appendix C.1). By contrast, a pure group of learning-aware agents cannot escape mutual defection, see Fig. 3C (right). This can be explained by the fact that the agents can no longer observe others learn, and must deal again with a non-stationary problem. The resulting gradients do not therefore contain information on the effects of unconditional defection on the future strategies of co-players, or that policies in the vein of tit-for-tat can shape co-players towards more cooperation.
@@
4.3 Explaining when and how cooperation arises with the LOLA algorithm
We next analyze the seminal Learning with Opponent-Learning Awareness (LOLA; Foerster et al., 2018a) algorithm. Briefly, LOLA assumes that co-players update their parameters with $M$ naive gradient steps, and estimates the total gradient through a look-ahead update:
@@ 4.3 Explaining when and how cooperation arises with the LOLA algorithm
with $\frac{\mathrm{d}}{\mathrm{d}\phi^{i}}$ the total derivative taking into account the effect of $\phi^{i}$ on the parameter updates $\Delta_{q}\phi^{-i}$, and $\frac{\partial}{\partial\phi^{-i}}$ the partial derivative. Note that Eq. 4 considers the LOLA-DICE update (Foerster et al., 2018b), an improved version of LOLA. In Appendix E, we show that Eq. 4 can be derived as a special case of COALA-PG. Note that LOLA-DICE estimates the policy gradient in Eq. 4 by explicitly backpropagating through the co-player’s parameter update using higher-order derivatives, whereas COALA-PG leads to a novel higher-order-derivative-free estimator of Eq. 4 (see Appendix E).
@@
Above, we showed that the two main ingredients for learning to cooperate under selfish objectives are (i) observe that one’s actions influence the future behavior of others, providing shaping gradients pulling away from defection towards extortion, and (ii), also play against other extortion agents immune to being shaped on the fast timescale, providing gradients pulling away from extortion towards cooperation. We then showed that both ingredients can be combined by training agents in a heterogeneous group containing both naive and learning-aware agents.
@@
Figure 4: (A) Performance of two agents trained by LOLA-DICE on the iterated prisoner’s dilemma with analytical gradients for various look-ahead steps (only the performance of the first agent is shown). (B) Performance of a randomly initialized naive learner trained against the fixed LOLA 20 look-aheads policy taken from the end of training of (A). (C) Same setting as (A), but with the naive gradient $\lambda\frac{\partial}{\partial\phi^{-i}}J^{i}(\phi,\phi^{-i})$ added to the LOLA-DICE update, with $\lambda$ a hyperparameter (c.f. Appendix C). Shaded regions indicate standard error computed over 64 seeds.
We can explain the emergent cooperation in LOLA by observing that LOLA also combines both ingredients, albeit differently from the heterogeneous group setting. The look-ahead rule (Eq. 4) computes gradients that shape naive learners performing $M$ naive gradient steps. Unique to LOLA however, these simulated naive learners are initialized with the parameters $\phi^{-i}$ of other LOLA agents. If the number of look-ahead steps is small, the updated naive learner parameters stay close to $\phi^{-i}$, mimicking playing against other extortion agents. This then results in emergent cooperation.
Fig. 4A confirms that LOLA-DICE with ground-truth gradients and with few look-ahead steps leads to cooperation on the iterated prisoner’s dilemma. However, as the number of look-ahead steps increases, the naive learner starts moving too far away from its $\phi^{-i}$ initialization, removing the second ingredient, thus leading to defection. In Fig. 4, we take the policy resulting from LOLA-training with many look-ahead steps, and train a new randomly initialized naive learner against this fixed LOLA policy. The results show that the LOLA policy extorts the naive learner into unfair cooperation, confirming that with many look-ahead steps, only a shaping incentive is present in the LOLA update, resulting in extortion policies. Hence, the low reward in Fig. 4A for LOLA agents with many look-ahead steps does not result from unconditional defection, but instead from both LOLA policies trying to extort the other one. Finally, we can improve the performance of LOLA with many look-ahead steps by explicitly introducing ingredient 2 by adding the partial gradient $\frac{\partial}{\partial\phi^{-i}}J^{i}(\phi,\phi^{-i})$ to Eq. 4, see Fig. 4C.
@@ 5 Experimental analysis of policy gradient implementations
The results presented in the previous sections were obtained by performing gradient ascent on analytical expected returns. This assumes knowledge of co-player parameters, and it is only possible on a restricted number of games which admit closed-form value functions. We now move to the general reinforcement learning setting, aiming at understanding (i) when meta-agents succeed in exploiting naive agents, and (ii) when cooperation is achieved among meta-agents.
@@ 5.1 Agents trained with COALA-PG master the iterated prisoner’s dilemma
Figure 5: Agents trained by COALA-PG play iterated prisoner’s dilemma. (A): When trained against naive agents only, COALA-PG-trained agents extort the latter and reach considerably higher reward than other baseline agents. The stars ($\star$) indicate overlapping curves of the corresponding color at that point (B): When analyzing the behavior of the agents within one meta-episode, we observe COALA-PG-trained agents shaping naive co-players, leading to low defection rate in the beginning, which is then exploited towards the end. M-FOS on the other hand defects from the beginning, achieving lower reward, thus failing to properly optimize the shaping problem. Batch-unaware COALA-PG performs identically to M-FOS and is therefore omitted. (C): Average performance of meta agents playing against other meta agents, when training a group of meta agents against a mixture of naive and other meta agents. Such agents trained with COALA-PG cooperate when playing against each other, but fail to do so when trained with baseline methods. When removing naive agents from the pool, meta agents also fail to cooperate, as predicted in Section 3. Shaded regions indicate standard deviation computed over 5 seeds.
We train a long-context sequence policy $\pi^{i}(a^{i,b}_{l}\mid h_{l}^{i,b};\phi^{i})$ with the COALA-PG rule to play the (finite) iterated prisoner’s dilemma, see Appendix B. We choose a Hawk recurrent neural network as the policy backbone (De et al., 2024). Hawk models achieve transformer-level performance at scale, but with time and memory costs that grow only linearly with sequence length. This allows processing efficiently the long history context $h_{l}^{i,b}$, which contains all actions played by the agents across episodes. Based on the results of the preceding section, we consider a mixed group setting, pitting COALA-PG-trained agents against naive learners as well as other equally capable learning-aware agents. Naive learners are equipped with the same policy architecture as the agents trained by COALA-PG, but their context is limited to the current inner game history $x_{t}^{i,b}$.
In Fig. 5, we see that COALA-PG reproduces the analytical game findings reported in the previous section: learning-aware agents cooperate with other learning-aware agents, and extort naive learners. Importantly, the identity of each agent is not revealed to a learning-aware agent, which must therefore infer in-context the strategy used by the player it faces. Likewise, we find that the LOLA-DICE (Eq. 4) estimator behaves as in the analytical game. This result complements previous experiments with LOLA on tabular policies (Foerster et al., 2018a, b), suggesting that there is a broad class of efficient learning-aware reinforcement learning rules that can reach cooperation with more complex context-dependent sequence models. We note that LOLA achieves this by explicitly differentiating through co-player updates, which requires access to their parameters and gives rise to higher-order derivatives. Our rule lifts these requirements, while maintaining learning efficiency.
@@
By contrast, the whole group falls into defection when training the exact same sequence model with the M-FOS rule, which weighs disproportionately future vs. current episode returns. We note that the experiments reported by Lu et al. (2022) were performed with a tabular policy and analytical inner game returns, for which cooperation could be achieved with the M-FOS rule. This shows how crucial the unbiased policy gradient property of COALA-PG is for co-player shaping by meta reinforcement learning to succeed in practice. The same failure to beat defection occurs when using a naive policy gradient ablation which does not take co-player batching into account. We refer to Appendix G for expressions for this baseline as well as the M-FOS rule. When training against M-FOS agents, our COALA-PG agents successfully shape M-FOS agents into cooperative behavior (c.f. Appendix H.3).
@@ 5.2 Agents trained with COALA-PG cooperate on a sequential social dilemma
Figure 6: Agents trained by COALA-PG against naive agents only successfully shape them in CleanUp-lite. (A) COALA-PG-trained agents better shape naive opponents compared to baselines, obtaining higher return. (B and C) Analyzing behavior within a single meta-episode after training reveals that COALA outperforms baselines and shapes naive agents, (i) exhibiting a lower cleaning discrepancy (absolute difference in average cleaning time between the two agents), and (ii) being less often zapped. Shaded regions indicate standard deviation computed over 5 seeds.
@@
Finally, we consider CleanUp-lite, a simplified two-player version of the CleanUp game, which is part of the Melting Pot suite of multi-agent environments (Agapiou et al., 2023). We briefly describe the game here, and provide additional details on Appendix B. On a high level, CleanUp-lite is a two-player game that models the social dilemma known as the tragedy of the commons (Hardin, 1968). A player receives rewards by picking up apples. Apples are spontaneously generated in an orchard, but the rate of generation is inversely proportional to the pollution level of a nearby river. Agents can spend time cleaning the river to reduce the pollution level, and thus increase the apple generation rate. In a single-player game, an agent would balance out cleaning and harvesting to maximize the return. In a multi-player setting however, this gives room for a “freerider" who never cleans and always harvests, letting the opponent clean instead. At any time point, agents can “zap" the opponent, which would result in the opponent being frozen for a number of time steps, unable to harvest or clean. In contrast to matrix games, this game is a sequential social dilemma (Leibo et al., 2017), where cooperation involves orchestrating multiple actions.
@@
As in the previous section, we model agent behavior through Hawk sequence policies, and compare COALA-PG to the same baseline methods as before. Naive agents here would learn too slowly if initialized from scratch, and are therefore handled differently, see Appendix B.2.2. In Figs. 6 and 7, we see that agents trained by COALA-PG reach significantly higher returns than previous model-free baselines, establishing a mutual cooperation protocol with other learning-aware agents, while exploiting naive ones. We further describe below the qualitative behavior found in the simulations.
@@
Exploitation of naive agents. COALA-PG-trained agents shape the behavior of naive ones to their advantage (c.f. Figure 6). Our behavioral analysis reveals two salient features. First, COALA-PG successfully shapes naive opponents to zap less often throughout the meta-episode. Less overall zapping means that agents can harvest more apples while the pollution level is low, thus increasing the overall reward. Second, COALA-PG successfully shapes naive co-players to clean significantly more often compared to the COALA-PG agent, resulting in a lower average pollution level and a higher average apple level (c.f. Figure 13 in App. H). Interestingly, the naive learners benefit from the shaping from COALA-PG agents, reaching a higher average reward compared to playing against other baselines (c.f. Figure 13).
@@
Figure 7: Agents trained with COALA-PG against a mixture of naive and other meta agents learn to cooperate in CleanUp-lite. (A) COALA-PG-trained agents obtain higher average reward than baseline agents when playing against each other. (B and C): COALA-PG leads to a more fair division of cleaning efforts and lower zapping rates. Shaded regions indicate standard deviation computed over 5 seeds.
@@
Learning-aware agents cooperate. We see similar trends when introducing other COALA-PG-trained agents in the game, see Fig. 7. Essentially, COALA-PG allows for higher apple production because of lower pollution, and lower zapping rate. We see that over training time the zapping rate goes down, and COALA-PG agents have a fairer division of cleaning time compared to baselines. Interestingly, the zapping rates averaged over the meta-episode are lower than in the pure shaping setting (i.e., with naive co-players only), indicating that learning-aware agents mutually shape each other to zap less.